Suppose i desk stock has a bela of 2 16, whereas Costco stock has a bela of 0.69. H the isk-free according to the CAPN? le est r te s 4% and the expected el m of the ma kel port lio s 10.0% what s he expecled return of a portfolio hat con sts of 60 Autodesk stock and 40% Costco stock. The expected return is--(Round to two decimal places.)
Answer #1

Expected return= risk-free rate +Beta*(market rate- risk-free rate )

Expected return of Autodesk=4+2.16*(10-4)=16.96%

Expected return of Costco=4+0.69*(10-4)=8.14%

Expected return of portfolio=Respective return*Respective investment weights

=(16.96*0.6)+(8.14*0.4)

=13.43%(Approx).